OpenStudy (anonymous):

can someone help please its solving equations 2/3x+1/2=7/2

OpenStudy (anonymous):

is it 2/3*x or 2/(3x)

OpenStudy (anonymous):

its 2/3*x

OpenStudy (anonymous):

multiply everything by 6

OpenStudy (anonymous):

ok thnx and then simplify right

OpenStudy (anonymous):

do that then i will tell you what to do

OpenStudy (anonymous):

ok i got 4x+3=21

OpenStudy (anonymous):

now move the 3 to the other side

OpenStudy (anonymous):

u mean subtrct from both sides or move it in front of the 4x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ook i got 4x+0=18| dw:1352770271240:dw|

OpenStudy (anonymous):

now divide by 4

OpenStudy (anonymous):

i got 4.5

OpenStudy (anonymous):

orif i do 4/18 it would be 22

OpenStudy (anonymous):

thnx very much

OpenStudy (anonymous):

18/4 = 9/2 = 4.5 depends on what the teacher wants